# A Computational Phase Function Approach for Obtaining $\alpha\alpha$ Wavefunctions

### v1

@@ Abstract
In this work, the phase function method (PFM) is applied for the first time to explicitly construct scattering wavefunctions for the $\alpha\alpha$ system using interaction parameters optimized from earlier phase shift analysis. While previous investigations employing PFM and related approaches have primarily focused on the reproduction of scattering phase shifts or cross sections, the present study advances the method to directly reconstruct physically meaningful radial scattering wavefunctions for the $\ell=0$, 2, and 4 partial waves without solving the Schrödinger equation. The computed wavefunctions exhibit well-behaved near-origin and asymptotic characteristics and show excellent agreement with the established resonating-group method results of Hiura et al.. This agreement confirms the numerical robustness and accuracy of the proposed framework. The present results establish PFM as an efficient, unified, and computationally attractive tool for scattering wavefunction reconstruction in cluster–cluster systems, opening new possibilities for its application to more complex nuclear scattering problems.
@@ 1 Introduction
An alternative and efficient framework is provided by the variable phase approach (VPA), also known as the phase function method (PFM), which reformulates the scattering problem in terms of first-order nonlinear differential equations for the radial phase shift and amplitude. Unlike traditional $S$-matrix or Jost-function methods [16, 17], PFM requires only the interaction potential as input and avoids explicit construction of the full scattering wave function at the outset. The method has been successfully applied to a variety of nuclear systems, including neutron–proton [22, 23], proton–proton [24], neutron–deuteron [25, 26, 27], and $\alpha$$\alpha$ scattering [28]. More recently, Laha and collaborators have demonstrated the versatility of PFM in studying nucleon–nucleon, nucleon–nucleus, and nucleus–nucleus interactions using different phenomenological potentials [18, 19, 20, 21].
The $\alpha$$\alpha$ system represents an especially suitable testing ground for the phase function framework. The $\alpha$ particle is characterized by zero spin and isospin, a large binding energy of 28.3 MeV, and a small root-mean-square radius of about 1.44 fm, allowing it to be treated as an elementary cluster. Consequently, the 8Be nucleus can be viewed as a weakly bound two-$\alpha$ system, and $\alpha$$\alpha$ scattering may be modeled effectively as a two-body problem governed by a local interaction potential. Since the pioneering experimental studies of Rutherford and Chadwick in 1927, a substantial body of experimental phase shift data has become available [4, 5, 6, 7, 8], enabling detailed theoretical investigations of the $\alpha$$\alpha$ interaction.
Historically, a variety of phenomenological potentials have been proposed to describe $\alpha$$\alpha$ scattering. Early studies by Häfner [9] and subsequent extensions by Nilson and collaborators [8] demonstrated the importance of angular-momentum dependence in reproducing experimental phase shifts. Later analysis using optical-model approaches and complex Woods–Saxon interactions revealed that no single $\ell$-independent potential can account for all partial waves over a wide energy range [6, 7]. Among simpler models, Buck et al. [12] employed a Gaussian nuclear interaction supplemented by a Coulomb term represented through an error function,
@@
The principal objective of this paper is to obtain a unified description of $\alpha$$\alpha$ scattering in terms of wave functions and scattering amplitudes within the phase function framework. By employing an inverse approach constrained by high-precision experimental phase shift data, the radial evolution of the phase shift, amplitude, and wave function is determined for relevant partial waves. This study aims to demonstrate that PFM, combined with physically motivated interaction models, provides deeper insight into the dynamics of the $\alpha$$\alpha$ system beyond conventional phase-shift analysis.
@@ 2.1 Morse Interaction Potential
The Morse potential exhibits several features that make it particularly suitable for modeling effective interactions between composite nuclear systems. It naturally incorporates a strong short-range repulsion, an attractive pocket at intermediate distances, and an exponentially decaying tail at large separations, all of which are characteristic of realistic nuclear forces. In addition, the time-independent Schrödinger equation with a Morse interaction is exactly solvable, leading to closed-form expressions for bound-state energies and wave functions. For scattering problems, the Morse potential admits analytical solutions for specific partial waves, most notably the $^{1}S_{0}$ state, a property not shared by many commonly used phenomenological interactions such as the Hulthén, Malfliet–Tjon, or Manning-Rosen potentials.
@@ 2.2 Coulomb Interaction
In addition to the short-range nuclear force, the long-range Coulomb repulsion between the two positively charged $\alpha$ particles is explicitly included. To account for the finite spatial extent of the $\alpha$ particle, the Coulomb interaction is represented using an error-function-modified form:
@@ 2.3 Phase Function Method (PFM)
Instead of solving the second-order differential equation (6) for the wave function, the phase function method (PFM), also known as the variable phase approach, reformulates the scattering problem in terms of a first-order nonlinear differential equation for the phase shift. This transformation relies on the mathematical equivalence between linear second-order differential equations and first-order Riccati equations, as originally developed by Calogero and Babikov.
@@ 2.3.1 Phase equations for $s$, $d$, and $g$ waves
2.4 Amplitude function A(r)
Within the phase function formalism, the radial amplitude $A_{\ell}(r)$ evolves with distance under the combined influence of the interaction potential and the running phase shift. The first-order differential equation governing the amplitude incorporates the local phase $\delta_{\ell}(k,r)$ and the Riccati–Bessel functions $\hat{j}_{\ell}(kr)$ and $\hat{\eta}_{\ell}(kr)$, thereby ensuring consistency with the partial-wave decomposition of the scattering solution. The general expression for the amplitude derivative is given by
$A_{\ell}^{\prime}(r)=$ $-\frac{A_{\ell}V(r)}{k}\left[\cos\left(\delta_{\ell}(k,r)\right)\hat{j}_{\ell}(kr)-\sin\left(\delta_{\ell}(k,r)\right)\hat{\eta}_{\ell}(kr)\right]$ (13)
$\times\left[\sin\left(\delta_{\ell}(k,r)\right)\hat{j}_{\ell}(kr)+\cos\left(\delta_{\ell}(k,r)\right)\hat{\eta}_{\ell}(kr)\right],$
Amplitude function equation for $\ell=0$ is given as
$A_{0}^{\prime}$ $=-\frac{A_{0}V(r)}{k\left(\frac{\hbar^{2}}{2\mu}\right)}\left[\cos\delta_{0}\sin(kr)+\sin\delta_{0}\cos(kr)\right]$
$\quad\times\left[\sin\delta_{0}\sin(kr)-\cos\delta_{0}\cos(kr)\right]$ (15)
Amplitude function equation for $\ell=2$ is given as
$A_{2}^{\prime}$ $=-\frac{A_{2}V(r)}{k\left(\frac{\hbar^{2}}{2\mu}\right)}\Bigg[\cos\delta_{2}\left(\left(\frac{3}{(kr)^{2}}-1\right)\sin(kr)-\frac{3}{(kr)}\cos(kr)\right)$
$\quad\times\Bigg(\sin\delta_{2}\left(\left(\frac{3}{(kr)^{2}}-1\right)\sin(kr)-\frac{3}{(kr)}\cos(kr)\right)$
$\qquad+\cos\delta_{2}\left(\left(-\frac{3}{(kr)^{2}}+1\right)\cos(kr)-\frac{3}{(kr)}\sin(kr)\right)\Bigg)\Bigg]$ (16)
Finally amplitude function equation for $\ell=4$ is given as
$A_{4}^{\prime}$ $=-\frac{A_{4}V(r)}{k\left(\frac{\hbar^{2}}{2\mu}\right)}\Bigg[\cos\delta_{4}\Bigg(\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\sin(kr)+\frac{10-105(kr)^{-2}}{kr}\cos(kr)\Bigg)$
$\quad-\sin\delta_{4}\Bigg(\left(-\frac{105}{(kr)^{2}}+10\right)\frac{\sin(kr)}{kr}+\left(-1+\frac{45}{(kr)^{2}}-\frac{105}{(kr)^{4}}\right)\cos(kr)\Bigg)\Bigg]$
$\quad\times\Bigg[\sin\delta_{4}\Bigg(\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\sin(kr)+\frac{10-105(kr)^{-2}}{kr}\cos(kr)\Bigg)$
$\qquad+\cos\delta_{4}\Bigg(\left(-\frac{105}{(kr)^{2}}+10\right)\frac{\sin(kr)}{kr}+\left(-1+\frac{45}{(kr)^{2}}-\frac{105}{(kr)^{4}}\right)\cos(kr)\Bigg)\Bigg]$ (17)
For the $s$-wave ($\ell=0$), the above equation simplifies due to the elementary forms of the spherical Bessel functions, leading to a comparatively compact expression for $A_{0}^{\prime}(r)$. The corresponding amplitude equation explicitly reflects the interplay between the potential strength and the running phase shift. A similar procedure is adopted for higher partial waves. In particular, the amplitude evolution for $\ell=2$ involves higher-order angular momentum terms arising from the centrifugal contribution, resulting in a more structured radial dependence. The amplitude equations for $\ell=0$ and $\ell=2$ are presented explicitly, while the corresponding expression for $\ell=4$ can be obtained analogously by employing the appropriate Riccati–Bessel functions.
The radial variation of the wavefunctions for $\ell=0$, $2$, and $4$ is illustrated in Fig. LABEL:fig2, highlighting the gradual stabilization of $A_{\ell}(r)$ at large distances where the interaction becomes negligible.
2.5 Wave function u(r)
@@
Once the amplitude and phase functions are determined, the radial wave function for each partial wave can be constructed directly. Within the phase function framework, the reduced radial wave function is expressed as
$u_{\ell}(r)=A_{\ell}(r)\left[\cos\left(\delta_{\ell}(k,r)\right)\hat{j}_{\ell}(kr)-\sin\left(\delta_{\ell}(k,r)\right)\hat{\eta}_{\ell}(kr)\right],$ (18)
which ensures the correct asymptotic behavior of the scattering solution.
Wavefunction equation for $\ell=0$ is given as
$u_{0}(r)$ $=A_{0}(r)\left[\cos\delta_{0}(r)\,\sin(kr)-\sin\delta_{0}(r)\,\cos(kr)\right]$ (20)
Wavefunction equation for $\ell=2$ is given as
$u_{2}(r)$ $=A_{2}(r)\Bigg[\cos\delta_{2}(r)\left(\left(\frac{3}{(kr)^{2}}-1\right)\sin(kr)-\frac{3}{(kr)}\cos(kr)\right)$
$\quad-\sin\delta_{2}(r)\left(\left(-\frac{3}{(kr)^{2}}+1\right)\cos(kr)+\frac{3}{(kr)}\sin(kr)\right)\Bigg]$ (21)
Finally wavefunction equation for $\ell=4$ is given as
$u_{4}(r)$ $=A_{4}(r)\Bigg[\cos\delta_{4}(r)\Bigg(\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\sin(kr)+\frac{10-105(kr)^{-2}}{kr}\cos(kr)\Bigg)$
$\quad-\sin\delta_{4}(r)\Bigg(\left(-\frac{105}{(kr)^{2}}+10\right)\frac{\sin(kr)}{kr}+\left(-1+\frac{45}{(kr)^{2}}-\frac{105}{(kr)^{4}}\right)\cos(kr)\Bigg)\Bigg]$ (22)
For $\ell=0$, the wave function reduces to a simple combination of sine and cosine functions modulated by the amplitude $A_{0}(r)$ and the running phase shift $\delta_{0}(r)$. In the case of $\ell=2$, the centrifugal barrier introduces additional radial structure through higher-order trigonometric terms, resulting in a more oscillatory behavior of the wave function at short distances. Explicit expressions for the $s$- and $d$-wave radial wave functions are provided, while the formulation for $\ell=4$ follows straightforwardly by substituting the corresponding spherical Bessel functions.
The radial dependence of the wave functions for $\ell=0$, $2$, and $4$ is displayed in Fig. 2, demonstrating the smooth transition from interaction-dominated behavior at small $r$ to asymptotic free-particle motion at larger separations.

### v2

# A Computational Phase Function Method for $\alpha-\alpha$ Scattering: Wavefunction Construction from Single and Two-Term Morse Potentials

@@ Abstract
[…]
For comparison, we adopt the interaction potential parameters reported by Sastri et al., who determined them using a two-term reference potential approach with genetic algorithm optimization to accurately reproduce the $\alpha\alpha$ scattering phase shifts. Without re-optimization, we construct the corresponding wavefunctions and find very good agreement with those obtained using our single-term Morse potential. The results also show excellent consistency with the resonating-group method calculations of Hiura et al..These findings demonstrate that PFM provides a numerically stable, efficient, and unified framework for scattering wavefunction construction in cluster–cluster systems.
@@ 1 Introduction
[…]
The $\alpha$$\alpha$ system represents an especially suitable testing ground for the phase function framework [22, 23]. The $\alpha$ particle is characterized by zero spin and isospin, a large binding energy of 28.3 MeV, and a small root mean square radius of about 1.44 fm, allowing it to be treated as an elementary cluster. Consequently, the 8Be nucleus can be viewed as a weakly bound two $\alpha$ system, and $\alpha$-$\alpha$ scattering may be modeled effectively as a two-body problem governed by a local interaction potential. Since the pioneering experimental studies of Rutherford and Chadwick in 1927, a substantial body of experimental phase shift data has become available [24, 25, 26, 27, 28], enabling detailed theoretical investigations of the $\alpha$$\alpha$ interaction.
Historically, a variety of phenomenological potentials have been proposed to describe $\alpha$$\alpha$ scattering. Early studies by Häfner [29] and subsequent extensions by Nilson and collaborators [28] demonstrated the importance of angular-momentum dependence in reproducing experimental phase shifts. Later analysis using optical-model approaches and complex Woods-Saxon interactions revealed that no single $\ell$-independent potential can account for all partial waves over a wide energy range [26, 27]. Among simpler models, Buck et al. [30] employed a Gaussian nuclear interaction supplemented by a Coulomb term represented through an error function,
@@
The principal objective of this paper is to obtain a unified description of $\alpha$$\alpha$ scattering in terms of wave functions and scattering amplitudes within the phase function framework. By employing an inverse approach constrained by high-precision experimental phase shift data, the radial evolution of the phase shift, amplitude, and wave function is determined for relevant partial waves. This study aims to demonstrate that PFM [35, 36], combined with physically motivated interaction models, provides deeper insight into the dynamics of the $\alpha$$\alpha$ system beyond conventional phase-shift analysis.
@@ 2.1 Morse Interaction Potential
The Morse potential exhibits several features that make it particularly suitable for modeling effective interactions between composite nuclear systems. It naturally incorporates a strong short-range repulsion, an attractive pocket at intermediate distances, and an exponentially decaying tail at large separations, all of which are characteristic of realistic nuclear forces. In addition, the time-independent Schrödinger equation with a Morse interaction is exactly solvable, leading to closed-form expressions for bound-state energies and wave functions. For scattering problems, the Morse potential admits analytical solutions for specific partial waves, most notably the $^{1}S_{0}$ state, a property not shared by many commonly used phenomenological interactions such as Manning-Rosen [19] potential, the Hulthén [21] potential or Malfliet-Tjon [38] potentials.
@@ 2.2 Coulomb Interaction
In addition to the short-range nuclear force, the long-range Coulomb repulsion between the two positively charged $\alpha$ particles is explicitly included. To account for the finite spatial extent of the $\alpha$ particle, the Coulomb interaction is represented using an error-function-modified form [31]:
@@
2.3 Two-Component Morse Reference Potential
The Reference Potential Approach (RPA), proposed by Selg [32, 33], provides a systematic procedure for treating one-dimensional quantum systems by selecting an analytically tractable reference interaction. Recently two-component Reference Potential Approach (RPA) was constructed from two smoothly connected Morse-type functions by Sastri et al. [34]. The composite potential is defined in a piecewise form as
$V_{\mathrm{RPA}}(r)=\begin{cases}V_{\mathrm{NC}}(r),&r\leq X,\\ V_{\mathrm{CL}}(r),&r\geq X,\end{cases}$ (6)
where $X$ is the matching (boundary) radius.
The inner-region potential is given by a standard Morse form,
$V_{\mathrm{NC}}(r)=V_{1}+D_{1}\left[e^{-2\alpha_{1}(r-r_{1})}-2e^{-\alpha_{1}(r-r_{1})}\right],$ (7)
and the outer-region potential is represented by a reversed Morse form,
$V_{\mathrm{CL}}(r)=V_{2}-D_{2}\left[e^{-2\alpha_{2}(r-r_{2})}-2e^{-\alpha_{2}(r-r_{2})}\right].$ (8)
Here,
  • $D_{k}$ $(k=1,2)$ denote the depth parameters,
  • $r_{k}$ are the equilibrium positions,
  • $\alpha_{k}$ control the range (shape) of each Morse component,
  • $V_{k}$ are constant offsets,
  • $X$ is the matching radius.
Continuity Conditions
To ensure smooth matching at $r=X$, both the potential and its first derivative must be continuous:
$V_{\mathrm{NC}}(X)=V_{\mathrm{CL}}(X),$ (9)
$\left.\frac{dV_{\mathrm{NC}}}{dr}\right|_{X}=\left.\frac{dV_{\mathrm{CL}}}{dr}\right|_{X}.$ (10)
Auxiliary Functions
For compact representation, define
$f_{1}$ $=e^{-2\alpha_{1}(X-r_{1})}-e^{-\alpha_{1}(X-r_{1})},$ (11)
$f_{2}$ $=e^{-2\alpha_{2}(X-r_{2})}-e^{-\alpha_{2}(X-r_{2})},$ (12)
$f_{3}$ $=e^{-2\alpha_{1}(X-r_{1})}-2e^{-\alpha_{1}(X-r_{1})},$ (13)
$f_{4}$ $=e^{-2\alpha_{2}(X-r_{2})}-2e^{-\alpha_{2}(X-r_{2})}.$ (14)
Determination of Depth Parameters
Imposing the continuity conditions yields the following expressions for $D_{1}$ and $D_{2}$:
$D_{1}=\frac{(V_{2}-V_{1})\alpha_{2}f_{2}}{\alpha_{2}f_{3}f_{2}-\alpha_{1}f_{4}f_{1}},$ (15)
$D_{2}=\frac{(V_{1}-V_{2})\alpha_{1}f_{1}}{\alpha_{2}f_{3}f_{2}-\alpha_{1}f_{4}f_{1}}.$ (16)
The resulting two-component Morse reference potential therefore depends on seven independent parameters:
$V_{1},V_{2},\alpha_{1},\alpha_{2},r_{1},r_{2},X,$
with $D_{1}$ and $D_{2}$ determined from the matching conditions. Sastri et al. determined the potential parameters using a genetic algorithm, which mitigates the risk of convergence to local minima. In the present work, we adopt the interaction potential constructed in Ref. [34] to compute the scattering phase shifts for the $\ell=0,2,$ and $4$ partial waves of $\alpha$$\alpha$ scattering. The results are systematically compared with those obtained from our single Morse potential formulation described in Sections 2.1 and 2.2.
@@ 2.4 Phase Function Method (PFM)
Instead of solving the second-order differential equation (17) for the wave function, the phase function method (PFM), also known as the variable phase approach, reformulates the scattering problem in terms of a first-order nonlinear differential equation for the phase shift. This transformation relies on the mathematical equivalence between linear second-order differential equations and first-order Riccati equations, as originally developed by Calogero [35] and Babikov [36].
@@ 2.4.1 Phase equations for $S$, $D$, and $G$ waves
2.5 Amplitude Function $A_{\ell}(r)$
Within the phase function framework, the scattering solution is expressed in terms of a running phase shift $\delta_{\ell}(k,r)$ and a radial amplitude function $A_{\ell}(r)$. While the phase equation governs the evolution of the local phase, the amplitude function determines the radial normalization of the scattering solution.
Starting from the phase–amplitude decomposition of the reduced radial wave function,
$u_{\ell}(r)=A_{\ell}(r)\Big[\cos\delta_{\ell}(k,r)\,\hat{j}_{\ell}(kr)-\sin\delta_{\ell}(k,r)\,\hat{\eta}_{\ell}(kr)\Big],$ (24)
where $\hat{j}_{\ell}(kr)$ and $\hat{\eta}_{\ell}(kr)$ denote the Riccati–Bessel and Riccati–Neumann functions, respectively, one obtains the following first-order differential equation for the amplitude function:
$A_{\ell}^{\prime}(r)=-\frac{V(r)}{k\left(\hbar^{2}/2\mu\right)}A_{\ell}(r)\Big[\cos\delta_{\ell}(k,r)\,\hat{j}_{\ell}(kr)-\sin\delta_{\ell}(k,r)\,\hat{\eta}_{\ell}(kr)\Big]\Big[\sin\delta_{\ell}(k,r)\,\hat{j}_{\ell}(kr)+\cos\delta_{\ell}(k,r)\,\hat{\eta}_{\ell}(kr)\Big].$ (25)
This expression is valid for all partial waves. The amplitude equation is integrated simultaneously with the phase equation from the origin outward, subject to the initial condition
$A_{\ell}(0)=1,$ (26)
which ensures regular normalization at short distances.
The evolution of $A_{\ell}(r)$ is driven by the interaction potential $V(r)$ and modulated by the local interference between regular and irregular free solutions. In the asymptotic region where $V(r)\rightarrow 0$, the derivative $A_{\ell}^{\prime}(r)\rightarrow 0$, and the amplitude approaches a constant value. This guarantees that the reconstructed wave function acquires the correct asymptotic scattering behavior.
For the specific partial waves considered in the present work, the required Riccati functions are:
S-wave ($\ell=0$)
$\hat{j}_{0}(kr)=\sin(kr),\qquad\hat{\eta}_{0}(kr)=-\cos(kr).$ (27)
D-wave ($\ell=2$)
$\hat{j}_{2}(kr)=\left(\frac{3}{(kr)^{2}}-1\right)\sin(kr)-\frac{3}{kr}\cos(kr),$ (28)
$\hat{\eta}_{2}(kr)=-\left(\frac{3}{(kr)^{2}}-1\right)\cos(kr)-\frac{3}{kr}\sin(kr).$ (29)
G-wave ($\ell=4$)
$\hat{j}_{4}(kr)=\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\sin(kr)+\left(\frac{10}{kr}-\frac{105}{(kr)^{3}}\right)\cos(kr),$ (30)
$\hat{\eta}_{4}(kr)=-\left(\frac{105}{(kr)^{4}}-\frac{45}{(kr)^{2}}+1\right)\cos(kr)+\left(\frac{10}{kr}-\frac{105}{(kr)^{3}}\right)\sin(kr).$ (31)
Substitution of these expressions into the general amplitude equation yields the explicit forms employed in the numerical calculations. Retaining the compact Riccati representation preserves clarity and avoids unnecessary algebraic expansion. These are provided seperately in Appendix section.
[…]
2.6 Wave Function Construction
@@
Once the phase function $\delta_{\ell}(k,r)$ and the amplitude function $A_{\ell}(r)$ are determined, the reduced radial scattering wave function can be reconstructed directly within the phase function framework. The solution is expressed in the phase–amplitude form as
$u_{\ell}(r)=A_{\ell}(r)\Big[\cos\delta_{\ell}(k,r)\,\hat{j}_{\ell}(kr)-\sin\delta_{\ell}(k,r)\,\hat{\eta}_{\ell}(kr)\Big],$ (32)
where $\hat{j}_{\ell}(kr)$ and $\hat{\eta}_{\ell}(kr)$ denote the Riccati–Bessel and Riccati–Neumann functions, respectively. This representation ensures that the reconstructed wave function satisfies the correct asymptotic scattering boundary condition.
In the asymptotic region $r\to\infty$, where the interaction potential $V(r)$ vanishes and the phase function saturates to its physical value $\delta_{\ell}(k)$, the wave function reduces to
$u_{\ell}(r)\longrightarrow A_{\ell}(\infty)\sin\!\left(kr-\frac{\ell\pi}{2}+\delta_{\ell}(k)\right),$ (33)
which is the standard form of the partial-wave scattering solution.
For the specific partial waves considered in this work, the explicit expressions are obtained by substituting the corresponding Riccati functions.
S-wave ($\ell=0$)
$u_{0}(r)=A_{0}(r)\sin\!\big(kr+\delta_{0}(k,r)\big).$ (34)
D-wave ($\ell=2$)
$u_{2}(r)=A_{2}(r)\Big[\cos\delta_{2}(k,r)\,\hat{j}_{2}(kr)-\sin\delta_{2}(k,r)\,\hat{\eta}_{2}(kr)\Big].$ (35)
G-wave ($\ell=4$)
$u_{4}(r)=A_{4}(r)\Big[\cos\delta_{4}(k,r)\,\hat{j}_{4}(kr)-\sin\delta_{4}(k,r)\,\hat{\eta}_{4}(kr)\Big].$ (36)
The radial behavior of $u_{\ell}(r)$ therefore reflects the combined evolution of the phase and amplitude functions. In the interaction region, deviations from free-particle oscillations encode the dynamical effects of the nuclear and Coulomb potentials, while at large distances the solution smoothly approaches the free asymptotic scattering form.
The phase function framework thus provides a direct and numerically stable procedure for constructing scattering wave functions without explicitly solving the second-order Schrödinger equation. The method yields simultaneous access to phase shifts, amplitudes, and radial wave functions within a unified formalism.